OpenStudy (anonymous):

solve 0.3x^2=1.8x+0.3

OpenStudy (anonymous):

help.

OpenStudy (anonymous):

start with \[.3x^2-1.8x -.3=0\]

OpenStudy (anonymous):

then multiply by ten to get \[3x^2-18x-3=0\] then perhaps the quadratic formula, or else factor if you can

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

kind of lost.

OpenStudy (turingtest):

did you forget the quadratic formula?

OpenStudy (anonymous):

no

OpenStudy (turingtest):

did you get the answer?

OpenStudy (anonymous):

i kind of got 2 square root of 5 +/-

OpenStudy (turingtest):

I'll check your work, one sec...

OpenStudy (turingtest):

no, that is wrong...

OpenStudy (turingtest):

the formula is\[x={-b \pm \sqrt{b^2-4ac}\over2a}\]for equations of the form \[ax^2+bx+c=0\] for you a=3 b=-18 c=-3 use that info and tell me what you get

OpenStudy (anonymous):

agree..but not exactly what am looking for the answer..

OpenStudy (turingtest):

o-0 are you saying you agree with the quadratic formula? I should hope so...

OpenStudy (anonymous):

my work is asking for the solution set??

OpenStudy (turingtest):

right, you will have two answers that make up that set as a result of the plus-or-minus part of the formula. so just plug in the numbers and simplify, what do you get?